A youth organization collected nickels and dimes for a charity drive. By the end of the 1st day drive the youth had collected 42.35, If there were 3 times as many dimes as nickels how many of each type of coin was collected?
Answer by TimothyLamb(4379)

10d + 5n = 4235
d = 3n
---
10(3n) + 5n = 4235
30n + 5n = 4235
35n = 4235
Answer:
n = 121
d = 363
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php